# Learning Confidence Bounds for Classification with Imbalanced Data

### v2

@@ 3.1 Training Data Bound
Let $X_{1},...,X_{k}$ be i.i.d. from a continuous distribution. Then $P(X_{a_{1}}<...<X_{a_{k}})=1/k!$ for any permutation $a_{1},...,a_{k}$ of $1,...,k$.
We denote $d_{n}=||\phi(x_{n})-\phi_{S}||$, where $\phi_{S}=\mathbb{E}[\phi(x)]$. In other words, $d_{n}$ is the Euclidean distance of a data point $x_{n}$ from its expected value in the projected space.
Given an i.i.d. training dataset $X_{t}=[x_{1},...,x_{N}]$ there exist $N!$ possible combinations for ordering the data. From this and using Theorem 1, the probability of a test point $x_{N+1}$ being further away from $\phi_{S}$ than any point in $X_{t}$ is
@@ 3.2 Concentration Inequalities
Let $S$ be an $N$ sample generated independently at random according to a distribution $P$ with support in a ball of radius $R$ in feature space, where $R=\sup_{n=1,...,N}||\phi(x_{n})||$ and $\bar{\phi}_{S}$ is the empirical mean of $S$. Then, with probability at least $1-\delta$ over the choice of $S$, we have
$||\bar{\phi}_{S}-\phi_{S}||\leq\frac{R}{\sqrt{N}}\Bigg{(}2+\sqrt{2\ln\frac{1}{\delta}}\Bigg{)}.$ (2)
From here, we can use the triangle inequality to bound the distance from a data point to its expected value, $d_{n}$, in terms of the distance from the empirical mean $\bar{d}_{n}=||\phi(x_{n})-\bar{\phi}_{S}||$:
@@ 4 Method
Let us assume we have access to a training dataset $(x_{1},y_{1}),\dots,(x_{n},y_{n})\in\mathcal{X}\times\mathcal{Y}$ with class labels $y\in\{\pm 1\}$, where samples $S_{1}$ and $S_{2}$ denote the negative and positive classes respectively. In order to separate the two classes, we can learn a classifier of the form $f(x)=sgn(\langle\phi(x),w\rangle+b)$ as described in Section 3. For example, in the case of a minimum-distance classifier, the weight vector would be given by $w=\bar{\phi}_{S_{1}}-\bar{\phi}_{S_{2}}$ and $b=\frac{1}{2}||\bar{\phi}_{S_{1}}-\bar{\phi}_{S_{2}}||$. Alternately, the scoring function $f(x)$ can be learnt by commonly used frameworks such as SVMs or Neural Networks.
@@
Assuming the half-spaces are non-overlapping implies that the sum of the supports of each individual class can be bounded by the difference in empirical means: $\hat{\bar{R}}_{1}+\hat{\bar{R}}_{2}\leq\hat{D}$, where $\hat{D}=||\langle\bar{\phi}_{S_{2}},w\rangle-\langle\bar{\phi}_{S_{1}},w\rangle||$ is the distance between the projections of the classes empirical means and $\hat{\bar{R}}_{i}=\sup_{x\in S_{i}}||\langle\phi(x),w\rangle-\langle\bar{\phi}_{S_{i}},w\rangle||$ is the empirical support of class $i$.
@@ 4.2 Relaxing the Constraints
The method proposed so far assumes linearly separable data in the projected space $\phi(x)$, with sufficient distance between the half-spaces such that Equation 7 can be satisfied. In practice this assumption will usually not hold. Therefore, we propose considering slack variables into the optimisation process.
@@
The optimisation of $L_{sv}$ is less straightforward than $L$. We iteratively set $\xi_{n}=1$ for class support instances, as these are the points directly contributing to $\hat{\bar{R}}_{i}$. Instances from each class are selected such that $\sum_{n\in\{1,...,N\}}\xi_{n}$ for each class matches the class imbalance present in the training set. The final method is described in pseudocode in Algorithm 1. We provide an alternative approach using continuous slack variables in Appendix A.
@@ Results
Baseline .$667\pm.02$$577\pm.04$$500\pm.05$
SMOTE [8] .$896\pm.03$$892\pm.03$$8$04
BW .$902\pm.02$$899\pm.02$$896\pm.03$
BMR [2] .$892\pm.04$$887\pm.05$$882\pm.06$
Thresh [37] .$892\pm.04$$887\pm.05$$882\pm.06$
Our Method .$\textbf{.911}\pm.01$$\textbf{.909}\pm.01$$\textbf{.914}\pm.01$
@@ 5.2 Class Imbalanced Datasets
Table 2 shows the details of the datasets. We scale all dataset features to $[-1,1]$. For the Hepatitis dataset we remove the features ‘Protime’, ‘Alkphosphate’ and ‘Albumin’ and 18 instances due to missing values. For the Heart disease dataset, we remove the features ‘ca’ and ‘thal’, as they contain missing values. To create an unbalanced dataset, we combine the heart disease types 3 and 4 as the positive class. For the MIMIC-III dataset [25], we perform ICU negative outcome prediction, using the data pre-processing outlined in [33].
@@ Results
Pima Diabetes Baseline .$540\pm.04$$181\pm.23$$137\pm.18$
SMOTE [8] .$595\pm.01$$495\pm.03$$396\pm.04$
BW .$585\pm.02$$476\pm.06$$373\pm.08$
BMR [2] .$661\pm.02$$628\pm.04$$575\pm.05$
Thresh [37] .$661\pm.02$$628\pm.04$$575\pm.05$
Our Method .$\textbf{.672}\pm.03$$\textbf{.651}\pm.05$$\textbf{.611}\pm.08$
Breast Cancer Baseline .$917\pm.02$$917\pm.02$$913\pm.02$
SMOTE [8] .$793\pm.09$$768\pm.12$$737\pm.14$
BW .$921\pm.02$$921\pm.02$$918\pm.02$
BMR [2] .$887\pm.08$$880\pm.09$$869\pm.11$
Thresh [37] .$935\pm.06$$932\pm.07$$929\pm.09$
Our Method .$\textbf{.943}\pm.02$$\textbf{.942}\pm.02$$\textbf{.946}\pm.02$
Hepatitis Baseline .$\textbf{.827}\pm.06$$623\pm.15$$485\pm.16$
SMOTE [8] .$804\pm.02$$299\pm.31$$224\pm.24$
BW .$768\pm.09$$681\pm.09$$499\pm.10$
BMR [2] .$804\pm.05$$656\pm.11$$493\pm.10$
Thresh [37] .$786\pm.08$$686\pm.09$$516\pm.09$
Our Method .$775\pm.09$$\textbf{.714}\pm.05$$\textbf{.535}\pm.08$
Heart Disease Baseline .$\textbf{.885}\pm.02$$784\pm.04$$718\pm.05$
SMOTE [8] .$844\pm.03$$700\pm.07$$602\pm.09$
BW .$842\pm.02$$781\pm.06$$664\pm.06$
BMR [2] .$4.8$04 .$819\pm.06$$704\pm.07$
Thresh [37] .$837\pm.05$$829\pm.03$$703\pm.06$
Our Method .$880\pm.02$$\textbf{.834}\pm.05$$\textbf{.742}\pm.06$
MIMIC ICU Baseline .$841\pm.01$$429\pm.03$$212\pm.02$
SMOTE [8] .$822\pm.01$$431\pm.02$$199\pm.01$
BW .$839\pm.01$$410\pm.02$$195\pm.02$
BMR [2] .$\textbf{.852}\pm.01$$405\pm.04$$2\pm.02$02
Thresh [37] .$844\pm.01$$421\pm.04$$208\pm.02$
Our Method .$802\pm.02$$\textbf{.477}\pm.04$$\textbf{.221}\pm.02$
@@ 6.1 Limitations
Appendix A Continuous Slack Variables
An alternative to the binary slack variables, $\xi_{n}$, presented in Sec. 4.2 is to use continuous values such that $\xi_{n}=0$ for $x_{i}$ that contribute towards $\hat{\bar{R}}_{i}$ and $N_{i}$, and $\xi_{n}=||x_{n}-\bar{\phi}_{S_{i}}||$ for $x_{i}$ which are allowed to exist outside of the supports of the data and do not contribute towards $\hat{\bar{R}}_{i}$ and $N_{i}$. This definition of slack variables is more similar to how they are defined in soft-margin SVMs [14], but adds additional complexity when compared to their binary counterpart.
@@ Appendix A Continuous Slack Variables
Pima Diabetes Baseline .$540\pm.04$$181\pm.23$$137\pm.18$
SMOTE [8] .$595\pm.01$$495\pm.03$$396\pm.04$
BW .$585\pm.02$$476\pm.06$$373\pm.08$
BMR [2] .$661\pm.02$$628\pm.04$$575\pm.05$
Thresh [37] .$661\pm.02$$628\pm.04$$575\pm.05$
Our Method .$\textbf{.672}\pm.03$$\textbf{.651}\pm.05$$\textbf{.611}\pm.08$
Breast Cancer Baseline .$917\pm.02$$917\pm.02$$913\pm.02$
SMOTE [8] .$793\pm.09$$768\pm.12$$737\pm.14$
BW .$921\pm.02$$921\pm.02$$918\pm.02$
BMR [2] .$887\pm.08$$880\pm.09$$869\pm.11$
Thresh [37] .$935\pm.06$$932\pm.07$$929\pm.09$
Our Method .$\textbf{.943}\pm.02$$\textbf{.941}\pm.02$$\textbf{.945}\pm.02$
Hepatitis Baseline .$\textbf{.827}\pm.06$$623\pm.15$$485\pm.16$
SMOTE [8] .$804\pm.02$$299\pm.31$$224\pm.24$
BW .$768\pm.09$$681\pm.09$$499\pm.10$
BMR [2] .$804\pm.05$$656\pm.11$$493\pm.10$
Thresh [37] .$786\pm.08$$686\pm.09$$516\pm.09$
Our Method .$775\pm.09$$\textbf{.714}\pm.06$$\textbf{.535}\pm.08$
Heart Disease Baseline .$\textbf{.885}\pm.02$$784\pm.04$$718\pm.05$
SMOTE [8] .$844\pm.03$$700\pm.07$$602\pm.09$
BW .$842\pm.02$$781\pm.06$$664\pm.06$
BMR [2] .$4.8$04 .$819\pm.06$$704\pm.07$
Thresh [37] .$837\pm.05$$829\pm.03$$703\pm.06$
Our Method .$878\pm.02$$\textbf{.835}\pm.05$$\textbf{.740}\pm.06$
MIMIC ICU Baseline .$841\pm.01$$429\pm.03$$212\pm.02$
SMOTE [8] .$822\pm.01$$431\pm.02$$199\pm.01$
BW .$839\pm.01$$410\pm.02$$195\pm.02$
BMR [2] .$\textbf{.852}\pm.01$$405\pm.04$$2\pm.02$02
Thresh [37] .$844\pm.01$$421\pm.04$$208\pm.02$
Our Method .$803\pm.02$$\textbf{.476}\pm.04$$\textbf{.220}\pm.02$

### v3

@@ 3.1 Training Data Bound
Let $X_{1},...,X_{k}$ be i.i.d. from a continuous distribution. Then $P(X_{a_{1}}<...<X_{a_{k}})=1/k!$ for any permutation $a_{1},...,a_{k}$ of 1,,k1,...,k.
[…]
Given an i.i.d. training dataset $X_{t}=\{x_{1},...,x_{N}\}$ there exist $N!$ possible combinations for ordering the data. From this and using Theorem 1, the probability of a test point $x_{N+1}$ being further away from $\phi_{S}$ than any point in $X_{t}$ is
@@
For a more detailed intuition behind deriving Eq. 1, refer to Appendix A.
@@ 3.2 Concentration Inequalities
[…]
$\bar{\phi}_{S}-\phi_{S}||\leq\frac{R}{\sqrt{N}}\Bigg(2+\sqrt{2\ln\frac{1}{\delta}}\Bigg).$ (2)
[…]
@@ 4 Method
[…]
@@
[…]
@@ 4.2 Relaxing the Constraints
The method proposed so far assumes linearly separable data in the projected space $\phi(x)$, with sufficient distance between the half-spaces such that Equation 7 can be satisfied. In practice this assumption will usually not hold. Therefore, we propose to consider slack variables into the optimisation process.
@@
The optimisation of $L_{sv}$ is less straightforward than $L$. We iteratively set $\xi_{n}=1$ for class support instances, as these are the points directly contributing to $\hat{\bar{R}}_{i}$. Instances from each class are selected such that $\sum_{n\in\{1,...,N\}}\xi_{n}$ for each class matches the class imbalance present in the training set. The final method is described in pseudocode in Algorithm 1. We provide an alternative approach using continuous slack variables in Appendix B.
@@ Results
Baseline $667\pm.02 .$$577\pm.04 .$$500\pm.05$
SMOTE [8] $896\pm.03 .$$92\pm.03 .88$$88\pm.04$
BW $902\pm.02 .$$899\pm.02 .$$896\pm.03$
BMR [2] $92\pm.04 .88$$7\pm.05 .88$$82\pm.06$
Thresh [37] $92\pm.04 .88$$7\pm.05 .88$$82\pm.06$
Our Method $\textbf{.911}\pm.01$ $\textbf{.909}\pm.01$ $\textbf{.914}\pm.01$
@@ 5.2 Class Imbalanced Datasets
Table 2 shows the details of the datasets. We scale all dataset features to [1,1][-1,1]. For the Hepatitis dataset we remove the features ‘Protime’, ‘Alkphosphate’ and ‘Albumin’ and 18 instances due to missing values. For the Heart disease dataset, we remove the features ‘ca’ and ‘thal’, as they contain missing values. To create an unbalanced dataset, we combine the heart disease types 3 and 4 as the positive class. For the MIMIC-III dataset [25], we perform ICU negative outcome prediction, using the data pre-processing outlined in [33].
@@ Results
Pima Diabetes
Baseline $540\pm.04 .$$181\pm.23 .$$137\pm.18$
SMOTE [8] $595\pm.01 .$$495\pm.03 .$$396\pm.04$
BW $585\pm.02 .$$476\pm.06 .$$373\pm.08$
BMR [2] $661\pm.02 .$$628\pm.04 .$$575\pm.05$
Thresh [37] $661\pm.02 .$$628\pm.04 .$$575\pm.05$
Our Method $\textbf{.672}\pm.03$ $\textbf{.651}\pm.05$ $\textbf{.611}\pm.08$
Breast Cancer
Baseline $917\pm.02 .$$917\pm.02 .$$913\pm.02$
SMOTE [8] $793\pm.09 .$$768\pm.12 .$$737\pm.14$
BW $921\pm.02 .$$921\pm.02 .$$918\pm.02$
BMR [2] $87\pm.08 .88$$80\pm.09$ $869\pm.11$
Thresh [37] $935\pm.06 .$$932\pm.07 .$$929\pm.09$
Our Method $\textbf{.943}\pm.02$ $\textbf{.942}\pm.02$ $\textbf{.946}\pm.02$
Hepatitis
Baseline $\textbf{.827}\pm.06$ $623\pm.15 .$$485\pm.16$
SMOTE [8] $804\pm.02 .$$99\pm.31 .22$$24\pm.24$
BW $768\pm.09 .$$681\pm.09 .$$499\pm.10$
BMR [2] $804\pm.05 .$$656\pm.11 .$$493\pm.10$
Thresh [37] $786\pm.08 .$$686\pm.09 .$$516\pm.09$
Our Method $775\pm.09 .$$\textbf{.714}\pm.05$ $\textbf{.535}\pm.08$
Heart Disease
Baseline $\textbf{.885}\pm.02$ $784\pm.04 .$$718\pm.05$
SMOTE [8] $844\pm.03 .$$700\pm.07 .$$602\pm.09$
BW $842\pm.02 .$$781\pm.06 .$$664\pm.06$
BMR [2] $848\pm.04 .$$819\pm.06 .$$704\pm.07$
Thresh [37] $837\pm.05 .$$829\pm.03 .$$703\pm.06$
Our Method $880\pm.02 .$$\textbf{.834}\pm.05$ $\textbf{.742}\pm.06$
MIMIC ICU
Baseline $841\pm.01 .$$429\pm.03 .$$212\pm.02$
SMOTE [8] $822\pm.01 .$$431\pm.02 .$$199\pm.01$
BW $839\pm.01 .$$410\pm.02 .$$195\pm.02$
BMR [2] $\textbf{.852}\pm.01$ $405\pm.04 .$$202\pm.02$
Thresh [37] $844\pm.01 .$$421\pm.04 .$$208\pm.02$
Our Method $802\pm.02 .$$\textbf{.477}\pm.04$ $\textbf{.221}\pm.02$
@@ 6.1 Limitations
Appendix A Training Data Bound
Here we elaborate on how to derive Eq. 1. First, we combine the training dataset $X_{t}=\{x_{1},...,x_{N}\}$ and test point, $x_{N+1}$, into one set $X=\{x_{1},...,x_{N},x_{N+1}\}$, where $|X|=N+1$.
Using the distance $d_{n}$ to order the set $X$, there are $(N+1)!$ possible combinations of ordering $X$. Out of these, there are $N!$ possible combinations when the distance to the test point distance, $d_{N+1}$, is greater than all other points in the set. Therefore we can define the probability of this occurring as
$\frac{N!}{(N+1)!}=\frac{N!}{N!(N+1)}=\frac{1}{(N+1)}.$
Appendix B Continuous Slack Variables
[…]
@@ Appendix B Continuous Slack Variables
Pima Diabetes
Baseline $540\pm.04 .$$181\pm.23 .$$137\pm.18$
SMOTE [8] $595\pm.01 .$$495\pm.03 .$$396\pm.04$
BW $585\pm.02 .$$476\pm.06 .$$373\pm.08$
BMR [2] $661\pm.02 .$$628\pm.04 .$$575\pm.05$
Thresh [37] $661\pm.02 .$$628\pm.04 .$$575\pm.05$
Our Method $\textbf{.672}\pm.03$ $\textbf{.651}\pm.05$ $\textbf{.611}\pm.08$
Breast Cancer
Baseline $917\pm.02 .$$917\pm.02 .$$913\pm.02$
SMOTE [8] $793\pm.09 .$$768\pm.12 .$$737\pm.14$
BW $921\pm.02 .$$921\pm.02 .$$918\pm.02$
BMR [2] $87\pm.08 .88$$80\pm.09$ $869\pm.11$
Thresh [37] $935\pm.06 .$$932\pm.07 .$$929\pm.09$
Our Method $\textbf{.943}\pm.02$ $\textbf{.941}\pm.02$ $\textbf{.945}\pm.02$
Hepatitis
Baseline $\textbf{.827}\pm.06$ $623\pm.15 .$$485\pm.16$
SMOTE [8] $804\pm.02 .$$99\pm.31 .22$$24\pm.24$
BW $768\pm.09 .$$681\pm.09 .$$499\pm.10$
BMR [2] $804\pm.05 .$$656\pm.11 .$$493\pm.10$
Thresh [37] $786\pm.08 .$$686\pm.09 .$$516\pm.09$
Our Method $775\pm.09 .$$\textbf{.714}\pm.06$ $\textbf{.535}\pm.08$
Heart Disease
Baseline $\textbf{.885}\pm.02$ $784\pm.04 .$$718\pm.05$
SMOTE [8] $844\pm.03 .$$700\pm.07 .$$602\pm.09$
BW $842\pm.02 .$$781\pm.06 .$$664\pm.06$
BMR [2] $848\pm.04 .$$819\pm.06 .$$704\pm.07$
Thresh [37] $837\pm.05 .$$829\pm.03 .$$703\pm.06$
Our Method $878\pm.02 .$$\textbf{.835}\pm.05$ $\textbf{.740}\pm.06$
MIMIC ICU
Baseline $841\pm.01 .$$429\pm.03 .$$212\pm.02$
SMOTE [8] $822\pm.01 .$$431\pm.02 .$$199\pm.01$
BW $839\pm.01 .$$410\pm.02 .$$195\pm.02$
BMR [2] $\textbf{.852}\pm.01$ $405\pm.04 .$$202\pm.02$
Thresh [37] $844\pm.01 .$$421\pm.04 .$$208\pm.02$
Our Method $803\pm.02 .$$\textbf{.476}\pm.04$ $\textbf{.220}\pm.02$